OpenStudy (anonymous):

(x to the 8th power +6)(x+8)?

OpenStudy (stormfire1):

\[(x^8 + 6)(x+8)\]\[x^9+8x^8+6x+48\]

OpenStudy (anonymous):

im so lost what do i do first?

OpenStudy (stormfire1):

Do you know how to FOIL?

OpenStudy (anonymous):

im not really that good at it , but i do know of it

OpenStudy (stormfire1):

Ok, well it's First Inner Outer Last First, multiply the first two terms of each expression: x^8 * x = x^9 Then multiple the inner terms: 6 * x = 6x Next, multiply the outer terms: x^8 * 8 = 8x^8 Now, the last term: 6 * 8 = 48 Put them all together and you'll get what I posted

OpenStudy (stormfire1):

err First Outer Inner Last I guess :) The order you do it doesn't matter...you just need to ensure you hit all four multiplications.

OpenStudy (anonymous):

so i would get 8x+48 right?

OpenStudy (stormfire1):

No, when you multiply two terms like (x + 1)(x+4), you will end up with four terms in the result...because you multiplied 4 times (FOIL). From there it may be possible to simplify but in this case it's not. The final answer is x^9 + 8x^8 + 6x + 48

OpenStudy (anonymous):

where did you get x to the 9 power

OpenStudy (stormfire1):

x^8 * x = x^9 (when you multiple exponents with the same base, you simply add the exponents...and x is the same as x^1)

OpenStudy (stormfire1):

This may help as well: http://www.freemathhelp.com/using-foil.html

OpenStudy (anonymous):

thanks

OpenStudy (stormfire1):

np...just remember that the point of FOIL is just to multiple each term in the first expression by each term in the second expression. FOIL is just a way to remember how to do it without missing anything

OpenStudy (anonymous):

i have another problem like that its (4x+2)(4x+2)

OpenStudy (stormfire1):

If you want, go ahead and post what you think you should do and I'll jump in if you do something wrong. This is pretty easy to learn once you do it a few times

OpenStudy (anonymous):

i would get 8x+16+8x+16 as the answer

OpenStudy (anonymous):

now i have to collect like terms right it would be 16x+16

OpenStudy (stormfire1):

you've got mistakes already :) do it step by step...what are the first two terms you need to multiply?

OpenStudy (anonymous):

the 4x and the 4x right i would times them though right?

OpenStudy (stormfire1):

Yes, and what would that = ?

OpenStudy (anonymous):

16x

OpenStudy (stormfire1):

no...you have to multiply the x's too. So it would be 16x^2

OpenStudy (stormfire1):

Now move on to the next term :)

OpenStudy (anonymous):

alright it would be 2*2 right?

OpenStudy (stormfire1):

yes...next one?

OpenStudy (anonymous):

that of course is 4 then you times it by that is where im lost

OpenStudy (anonymous):

do i multiply 4x*2 =8 then 4x*2= 8 again then just add them

OpenStudy (stormfire1):

exactly...since they are both like terms of 8x, the result will be 16x

OpenStudy (stormfire1):

Now that you've done that, you'll end up with 16x^2 + 16x + 4

OpenStudy (stormfire1):

which is the answer

OpenStudy (stormfire1):

Of course the order of the terms in the final answer doesn't matter but it's generally expressed from the highest exponent to the lowest...which is why I ordered it that way.

OpenStudy (anonymous):

where did i get 16 to the 2nd power

OpenStudy (stormfire1):

remember when we did 4x *4x? That = 16x^2

OpenStudy (anonymous):

oh alright now i got it

OpenStudy (stormfire1):

The essence of this is just multiplying every term in each set of parentheses with every term in the other sets. If you had 3 sets of parentheses (like (x+1)(2x+4)(3x-1), you'd have to do this for each term in each set of parenthesis...which gets pretty ugly.

OpenStudy (anonymous):

thank you

OpenStudy (stormfire1):

Luckily, right now you're only working with 2 factors (sets of parentheses)

OpenStudy (stormfire1):

no, problem...good luck :)